# The $\ell$-test: leveraging sparsity in the Gaussian linear model for improved inference

### v1

@@ 5 Experiments
In this section, we perform experiments to evaluate the performance of the $\ell$-test and its corresponding confidence intervals and post-selection procedures. For all simulations except in Section 5.2 (where we study the robustness of the $\ell$-test to deviations from model (1.1)), we use a linear model (1.1) with $k$ out of the $d$ elements of $\bm{\beta}$ chosen uniformly without replacement and set to $A$ or $-A$ with equal probability, and all the other remaining entries set to 0. We perform inference on one randomly chosen signal coefficient, $\beta_{j}$. The rows of $\bm{X}$ are drawn i.i.d. from $\mathcal{N}_{d}(\bm{0},\bm{\Sigma})$ and then the columns are normalized. We will specify the values of $n,d,k,A,\bm{\Sigma}$, and $\sigma$ for each of the simulation settings we consider.
@@ 5.1 Power of the $\ell$-test
We compare the power of the following three tests: The $\ell$-test, the two-sided $t$-test, and the one-sided $t$-test in the direction of the true sign of $\beta_{j}$, under simulation settings studying the effect of varying the amplitude of the signal variables, the number of signal variables, and the inter-variable correlation. Note there is no need to compare Type I error rates, since all three methods have guaranteed exactly nominal Type I error (as long as model (1.1) is well-specified, which it is in this subsection). The results are reported in Figure 2 (see the caption for further details of the simulations). A simulation considering the case where $d$ is closer to $n$ is provided in Appendix F.1; the agreement between the $\ell$-test and the one-sided $t$-test becomes even stronger in this case.
@@ 5.2 Robustness of the $\ell$-test
Figure 3: We vary the sample size ($n$) on the x-axis and set $d=n/2$. For the top-left, we draw $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}t_{2}$ (which does not have a finite second moment), for the top-right, $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathrm{Exp}(1)$, standardized with its theoretical mean and standard deviation, and for the bottom-left, $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,1)$ if mean of the $i^{\mathrm{th}}$ row of $\bm{X}$ is less than the median of the row-means, while, $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,8)$, otherwise. For bottom-right, we generate $y_{i}\sim\mathcal{N}\left((\bm{X}_{i}^{4})^{T}\bm{\beta},\bm{I}\right)$, where, $(\bm{X}_{i}^{4})_{j}=X_{ij}^{4}$. All these settings use the nominal size of $5\%$ and the error bars represent plus or minus two units of standard error.
One thing that makes the $t$-test remarkable and so useful in practice is its robustness, even in relatively small samples, to violations of model (1.1). To evaluate the $\ell$-test’s robustness, we fix $d=n/2,k=1,\bm{\Sigma}=\bm{I}$, $A=3.3$, and test on a null index $\beta_{j}=0$. Figure 3 shows Type I error results for the $\ell$-test and $t$-test for four types of model violation: heavy-tailed errors, skewed errors, heteroskedastic errors, and model non-linearity (the figure caption gives exact specifications of each of the model violations) for a range of small sample sizes. In Appendix F.2 we present further simulations of these same four types of model violations, but Figure 3 shows the most extreme example of each of the four. Despite substantial deviations from (1.1), the $\ell$-test remains quite robust: like the $t$-test, it only makes even noticeable Type I error violations in the case of (very subtantial) heteroskedasticity, but even in this case (which is a well-known source of non-robustness for the $t$-test), the Type I error violation of the $\ell$-test remains close to that of the $t$-test and both are just a few percentage points above nominal.
@@ 5.3 Confidence Intervals for $\beta_{j}$
We now perform simulations to compare the $\ell$-test confidence interval with the usual (two-sided) $t$-test confidence interval, as well as the interval obtained by inverting the one-sided $t$-test in the direction of the true sign of the alternative for every alternative $\beta_{j}=\gamma$. In Appendix G, we discuss an explicit characterization of this last interval, however a point we highlight here is that this is an oracle procedure (even more so than the one-sided $t$-test we compare to in Section 5.1) and can only be constructed if we know the exact true value of the coefficient, not just its sign. To obtain the $\ell$-test confidence interval, we choose a grid of candidate values for $\beta_{j}$ and report the coverage and length of the smallest interval that strictly encloses the rejected values of $\beta_{j}$ from both the ends. Note that this interval will always have length and coverage at least as large as the true $\ell$-test confidence interval (which we can never obtain exactly from a finite grid of $\beta_{j}$ values). We use brute force calculations using the highly optimized functions available in the package glmnet (Friedman et al., 2010) in R instead of our theoretically efficient algorithm in Appendix D for $\ell$-test inversion (and defer the task of designing an optimized implementation for it to future research).
For our simulations we consider the exact settings as in Figure 2 and summarize the results in Figure 4. We only report the lengths of all the intervals for a more compact presentation whereas the full results with empirical converges are reported in Appendix F.3 (the coverage is always extremely close to the nominal 95%). As expected, we see similar trends as in Figure 2, with the $\ell$-test confidence intervals being close to the oracle one-sided $t$-test intervals, consistently across amplitudes, in sparse settings. In the left and right plots, the $\ell$-test confidence intervals are consistently about $12\%$ shorter than their $t$-test based counterparts. Perhaps surprisingly, the center plot shows that the $\ell$-test’s benefit in confidence interval width over the $t$-test remains up until about 80% nonzero entries in the coefficient vector, which is a larger outperformance range of sparsity than in Figure 2. Similar to the power results, we see only a small detriment to using the $\ell$-test confidence interval in the densest setting, relative to its benefit in the sparsest setting. As with the power simulations, we also consider a setting with $d$ closer to $n$ in Appendix F.3, and again find this further narrows the gap between the $\ell$-test and one-sided $t$-test procedures.
@@ 5.4 Post-selection $\ell$-test inference
In this section, we perform simulations to empirically evaluate the performance of the adjusted $\ell$-test confidence intervals in Section 4 for post-selection inference in the linear model under LASSO selection. In particular, we will compare $\underline{\hat{C}}_{j}^{\lambda}$ and $\underline{\hat{C}}_{j}^{*\lambda}$, where for both of these methods we invert the respective adjusted $\ell$-tests on a grid of values using the same strategy as in Section 5.3, along with the conditional confidence interval procedure of Liu et al. (2018). Figure 5 shows the effect of varying coefficient amplitude on the length and coverage of the intervals (conditional on the LASSO selecting the coefficient) for the same setting as the left panels of Figures 2. We see there is practically no difference between the performance of $\underline{\hat{C}}_{j}^{\lambda}$ and $\underline{\hat{C}}_{j}^{*\lambda}$ (see figure caption for how $\lambda$ was chosen) and both are consistently shorter than the method in Liu et al. (2018), reflecting again the benefits of the $\ell$-test under sparsity but now conditional on LASSO selection. In Appendix F.4, we show results for an additional setting where $d$ is closer to $n$.
@@ 5.5 Analysis of the HIV drug resistance data
The HIV drug resistant data (Rhee et al., 2006) consists of 16 different regressions, each containing data on a set of genetic mutations (the covariates) and a score measuring resistance to an HIV drug (the response). We follow the same pre-processing step suggested in Barber and Candès (2015), resulting in 16 regressions with $n$ ranging between 328 and 842 and $d$ ranging between 147 and 313. Running the $t$-test on all covariates across all data sets results in an average power of 16.7%, while for the $\ell$-test it is 18.6% (this represents an 11% improvement), showing that the $\ell$-test’s theoretical benefits under sparsity provide genuine power gains in real data sets. Similarly, the $t$-test confidence intervals have an average width of 3.85 while the $\ell$-test confidence intervals’ is 3.54, representing about an 8% improvement. We also note that previous works studying the same data have aggregated discovered mutations to the gene level, and if we do this, the comparison remains similar: the $t$-test’s average power is 31.8% discovered genes while the $\ell$-test’s is 36.4% (a 14% improvement).
@@ 6 Discussion
The $\ell$-test leverages sparsity by using the LASSO coefficient estimate $|\hat{\beta}^{\lambda}_{j}|$ as its test statistic and can achieve power close to that of the one-sided $t$-test without any knowledge about the true sign of the coefficient. The $\ell$-test can be inverted to confidence intervals that are over 10% shorter than $t$-test intervals under sparsity, and the $\ell$-test and confidence intervals can also be adjusted for LASSO selection. A number of questions remain for future work:
  1. 1.
    A recentered $u_{1}$-based test. Section 2.2 argued that the $\ell$-test’s power gains under sparsity are derived from $\hat{m}_{j}$ tending to take the opposite sign of $\beta_{j}$, as this results in the $\ell$-test p-value (which uses the test statistic $|\hat{\beta}_{j}^{\lambda}|$) putting most of its weight on the “correct” tail of $\hat{\beta}_{j}^{\lambda}$. In fact, due to the increasing relationship between $u_{1}$ and $\hat{\beta}_{j}^{\lambda}$ established in Theorem 2.1, it is easy to see that a similar phenomenon occurs if we use $|u_{1}-\hat{m}_{j}|$ as the test statistic (note this is the same test statistic used to break ties in the $\ell$-test in Section 2.3), and indeed, although doing so produces a test that is numerically distinct from the $\ell$-test, the two tests perform very similarly. We presented the $\ell$-test in its current form because we expect it to be more interpretable for most (especially non-statistician) users to have its test statistic be a natural estimator of $\beta_{j}$ under sparsity, as well as because of the easy extension to post-selection inference, but recognizing that the test statistic $|u_{1}-\hat{m}_{j}|$ produces a very similar test may be helpful in generalizing the $\ell$-test idea or for theoretical study such as power analysis.
  2. 2.
    Extension to Gaussian means and, asymptotically, to general parametric models. A Gaussian means problem, wherein $\bm{w}\sim\mathcal{N}(\bm{\beta},\bm{\Sigma})$ is observed for some known positive-definite $\bm{\Sigma}$ (but $\bm{\beta}$ unknown), can always be converted to linear regression by taking $\bm{X}=\bm{\Sigma}^{-1/2}$ and $\bm{y}=\bm{\Sigma}^{-1/2}\bm{w}$, so that $\bm{y}\sim\mathcal{N}(\bm{X}\bm{\beta},\bm{I}_{n})$. The key point is that the $\ell$-test [^3] When $\sigma^{2}$ is known, the $\ell$-test can easily use this information for a slight improvement by shrinking $\bm{S}^{(j)}$ to just its first element $\tilde{\bm{S}}^{(j)}:=\bm{X}_{-j}^{T}\bm{y}$ and replacing Equation (2.2) with the statement that $\hat{\sigma}_{j}\bm{u}\mid\tilde{\bm{S}}^{(j)}\sim\mathcal{N}(0,\sigma^{2}\bm{I}_{n-d+1})$. Theorem 2.1 remains unchanged, and in particular Equation (2.5) equates the event $\{\hat{\beta}_{j}^{\lambda}\leq b\}$ with $\left\{\hat{\sigma}_{j}u_{1}\leq\hat{\sigma}_{j}\Lambda_{j}(b,\mathrm{sign}(b))\right\}$, where $\hat{\sigma}_{j}\Lambda_{j}(b,\mathrm{sign}(b))$ is easily seen to be a function only of $\tilde{\bm{S}}^{(j)}$ as required for straightforward evaluation of the probability of this event via the Gaussian CDF. can be directly applied to any Gaussian means problem with known covariance matrix, with the analogous expectation that, when $\bm{\beta}$ is sparse, the $\ell$-test will outperform the standard $z$-test as long as $\bm{\Sigma}$ has sufficient off-diagonal entries (recall from Section 2.2 that the $\ell$-test’s power gain under sparsity derives from the sign-guessing ability of $\hat{m}_{j}$, which in turn relies on non-orthogonality of the columns of $\bm{X}$). This opens up many possible further applications of the $\ell$-test, including to any multivariate parametric model that admits an asymptotically multivariate Gaussian estimator and a consistent estimator for its covariance matrix, which includes both classical low-dimensional maximum likelihood or M-estimator asymptotics (see (Li and Fithian, 2021, Theorem 4) for an asymptotic result like this for knockoffs) as well as certain proportional asymptotic regimes with $n\propto d$ (Sur and Candès, 2019, Zhao et al., 2022).
  3. 3.
    Leveraging structure other than sparsity. The $\ell$-test leverages sparsity to improve upon the $t$-test, but it would be helpful to have analogous methods to use in settings that may not be sparse, but are believed to satisfy other forms of structure. For instance, if $\bm{\beta}$ is dense but smooth (i.e., it has small total variation), a test based on the fused LASSO (Tibshirani et al., 2005) may be more appropriate. The FAB method (Hoff, 2022) mentioned in Section 1.3 achieves a similar goal as the $\ell$-test for other forms of prior information and it may also be interesting to understand the relationship between the two.
@@ Appendix A Characterization of the $\ell$-distribution under $H_{j}(\gamma):\beta_{j}=\gamma$
Theorem A.1.
Consider the linear model defined in Theorem 2.1, fix $\gamma\in\mathbb{R}$ and define $\hat{\sigma}_{j}(\gamma)=\|(\bm{I}-\bm{P}_{-j})(\bm{y}-\gamma\bm{X}_{j})\|$. Then,
  1. 1.
    $\bm{S}^{(j)}(\gamma)=(\bm{X}_{-j}^{T}\bm{y},\|(\bm{I}-\bm{P}_{-j})(\bm{y}-\gamma\bm{X}_{j})\|)$ is sufficient under $H_{j}(\gamma)$. Furthermore, fixing an orthogonal matrix $\bm{V}$ for the column-space of $\bm{X}_{-j}$ as in Lemma 2.1, there exists a unique vector $\bm{u}^{\gamma}\in\mathbb{S}^{n-d}$ such that $\bm{y}=\hat{\bm{y}}_{j}+\gamma(\bm{I}-\bm{P}_{-j})\bm{X}_{j}+\hat{\sigma}_{j}(\gamma)\bm{V}\bm{u}^{\gamma}$ and
@@
Note that, as discussed in Section 4, the proof of item 1 of the above theorem directly follows from Lemma 2.1 applied to $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$. Analogous to the proof of Theorem 2.1 using Lemma 2.1, one can derive item 2 of Theorem A.1 from its item 1, by defining $\bm{z}:=\bm{y}-\gamma\bm{X}_{j}$ and $\bm{\delta}$ by $\bm{\delta}_{-j}=\bm{\beta}_{-j}$ and $\delta_{j}=\beta_{j}-\gamma$ and copying exactly the same proof as in Appendix C.3 but by replacing $(\bm{y},\bm{\beta},\hat{\sigma}_{j})$ with $(\bm{z},\bm{\delta},\hat{\sigma}_{j}(\gamma))$ and then substituting back for $\bm{\delta}$ and $\bm{z}$ at the end. We thus skip an explicit proof of Theorem A.1.
Here we would also like to draw attention to the fact that the function $f_{\bm{S}^{(j)}(\gamma)}^{\gamma}$, for any $\gamma\in\mathbb{R}$, is defined for any real number $u$ and not necessarily restricted to $[0,1]$ and also that $\Lambda_{j}(b,\mathrm{sign}(b))$ can take values outside the $[-1,1]$ range. In fact if the values exceed this range, we can often draw conclusive insights about the behavior of the LASSO estimate $\hat{\beta}_{j}^{\lambda}$. For example, $\Lambda(0,\pm 1)<-1$ implies that $u_{1}^{\gamma}>\Lambda(0,1)$ and hence that for any $\bm{y}$ generated using the condition in Theorem A.1 for a $\bm{u}^{\gamma}\in\mathbb{S}^{n-d}$, the resultant LASSO estimator of the $j^{\mathrm{th}}$ coefficient, $\hat{\beta}_{j}^{\lambda}$, will always be positive. This can happen in situations as we next described in Section B and can in-fact, result in the $\ell$-test producing exactly the one-sided $t$-test p-value.

### v2

@@ 6 Experiments
In this section, we perform experiments to evaluate the performance of the $\ell$-test and its corresponding confidence intervals and post-selection procedures. For all simulations except in Section 6.2 (where we study the robustness of the $\ell$-test to deviations from model (1.1)), we use a linear model (1.1) with $k$ out of the $d$ elements of $\bm{\beta}$ chosen uniformly without replacement and set to $A$ or A-A with equal probability, and all the other remaining entries set to 0. We perform inference on one randomly chosen signal coefficient, $\beta_{j}$. The rows of $\bm{X}$ are drawn i.i.d. from $\mathcal{N}_{d}(\bm{0},\bm{\Sigma})$ and then the columns are normalized. We will specify the values of $n,d,k,A,\bm{\Sigma}$, and $\sigma$ for each of the simulation settings we consider.
@@ 6.1 Power of the $\ell$-test
We compare the power of the following three tests: The $\ell$-test, the two-sided $t$-test, and the one-sided $t$-test in the direction of the true sign of $\beta_{j}$, under simulation settings studying the effect of varying the amplitude of the signal variables, the number of signal variables, and the inter-variable correlation. Note there is no need to compare Type I error rates, since all three methods have guaranteed exactly nominal Type I error (as long as model (1.1) is well-specified, which it is in this subsection). The results are reported in Figure 2 (see the caption for further details of the simulations). A simulation considering the case where $d$ is closer to $n$ is provided in Appendix H.1; the agreement between the $\ell$-test and the one-sided $t$-test becomes even stronger in this case.
@@ 6.2 Robustness of the $\ell$-test
Figure 3: We fix $\rho=0$ and vary the sample size ($n$) on the x-axis and set $d=n/2$. For the top-left, we draw $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}t_{2}$ (which does not have a finite second moment), for the top-right, $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}\mathrm{Exp}(1)$, standardized with its theoretical mean and standard deviation, and for the bottom-left, $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,1)$ if mean of the $i^{\mathrm{th}}$ row of $\bm{X}$ is less than 0, while $\epsilon_{i}\stackrel{{\scriptstyle i.i.d.}}{{\sim}}N(0,8)$ otherwise. For bottom-right, we generate $y_{i}\sim\mathcal{N}\left((\bm{X}_{i}^{4})^{T}\bm{\beta},\bm{I}\right)$, where, $(\bm{X}_{i}^{4})_{j}=X_{ij}^{4}$. All these settings use the nominal size of $5\%$ and the error bars represent plus or minus two units of standard error.
One thing that makes the $t$-test remarkable and so useful in practice is its robustness, even in relatively small samples, to violations of model (1.1). To evaluate the $\ell$-test’s robustness, we fix $d=n/2,k=1,A=3.3$, $\bm{\Sigma}$ to have a Toeplitz structure with $\Sigma_{i,j}=\rho^{|i-j|}$ and test on a null index $\beta_{j}=0$. Figure 3 shows Type I error results for the $\ell$-test and $t$-test for four types of model violation, when we choose $\bm{\Sigma}=\bm{I}$ (that is, $\rho=0$): heavy-tailed errors, skewed errors, heteroskedastic errors, and model non-linearity (the figure caption gives exact specifications of each of the model violations) for a range of small sample sizes. In Appendix H.2 we present further simulations of these same four types of model violations, but Figure 3 shows the most extreme example of each of the four. Appendix H.2 also presents the counterparts of these settings that includes correlation between the variables by setting $\rho=0.5$. Despite substantial deviations from (1.1), the $\ell$-test, like the $t$-test, remains quite robust—their Type I error violations remain close and both are at most just a few percentage points above nominal.
@@ 6.3 Confidence Intervals for $\beta_{j}$
We now perform simulations to compare the $\ell$-test confidence interval with the usual (two-sided) $t$-test confidence interval, as well as the interval obtained by inverting the one-sided $t$-test in the direction of the true sign of the alternative for every alternative $\beta_{j}=\gamma$. In Appendix I, we discuss an explicit characterization of this last interval, however a point we highlight here is that this is an oracle procedure (even more so than the one-sided $t$-test we compare to in Section 6.1) and can only be constructed if we know the exact true value of the coefficient, not just its sign. To obtain the $\ell$-test confidence interval, we choose a grid of candidate values for $\beta_{j}$ and report the coverage and length of the smallest interval that strictly encloses the rejected values of $\beta_{j}$ from both the ends. Note that this interval will always have length and coverage at least as large as the true $\ell$-test confidence interval (which we can never obtain exactly from a finite grid of $\beta_{j}$ values). We use brute force calculations using the highly optimized functions available in the package glmnet (glmnet1) in R instead of our theoretically efficient algorithm in Appendix F for $\ell$-test inversion (and defer the task of designing an optimized implementation for it to future research).
For our simulations we consider the exact settings as in Figure 2 and summarize the results in Figure 4. We only report the lengths of all the intervals for a more compact presentation whereas the full results with empirical converges are reported in Appendix H.3 (the coverage is always extremely close to the nominal 95%). As expected, we see similar trends as in Figure 2, with the $\ell$-test confidence intervals being close to the oracle one-sided $t$-test intervals, consistently across amplitudes, in sparse settings. In the left and right plots, the $\ell$-test confidence intervals are consistently about $12\%$ shorter than their $t$-test based counterparts. Perhaps surprisingly, the center plot shows that the $\ell$-test’s benefit in confidence interval width over the $t$-test remains up until about 80% nonzero entries in the coefficient vector, which is a larger outperformance range of sparsity than in Figure 2. Similar to the power results, we see only a small detriment to using the $\ell$-test confidence interval in the densest setting, relative to its benefit in the sparsest setting. As with the power simulations, we also consider a setting with $d$ closer to $n$ in Appendix H.3, and again find this further narrows the gap between the $\ell$-test and one-sided $t$-test procedures.
@@ 6.4 Post-selection $\ell$-test inference
In this section, we perform simulations to empirically evaluate the performance of the adjusted $\ell$-test confidence intervals in Section 5 for post-selection inference in the linear model under LASSO selection. In particular, we will compare $\underline{\hat{C}}_{j}^{\lambda}$ and $\underline{\hat{C}}_{j}^{*\lambda}$, where for both of these methods we invert the respective adjusted $\ell$-tests on a grid of values using the same strategy as in Section 6.3, along with the conditional confidence interval procedure of liu2018powerful. Figure 5 shows the effect of varying coefficient amplitude on the length and coverage of the intervals (conditional on the LASSO selecting the coefficient) for the same setting as the left panels of Figure 2. We see there is practically no difference between the performance of $\underline{\hat{C}}_{j}^{\lambda}$ and $\underline{\hat{C}}_{j}^{*\lambda}$ (see figure caption for how $\lambda$ was chosen) and both are consistently shorter than the method in liu2018powerful, reflecting again the benefits of the $\ell$-test under sparsity but now conditional on LASSO selection. In Appendix H.4, we show results for an additional setting where $d$ is closer to $n$.
@@ 6.5 Analysis of the HIV drug resistance data
The HIV drug resistance data (hivdata) consists of 16 different regressions, each containing data on a set of genetic mutations (the covariates) and a score measuring resistance to an HIV drug (the response). We follow the same pre-processing step suggested in knockoff, resulting in 16 regressions with $n$ ranging between 328 and 842 and $d$ ranging between 147 and 313. Running the $t$-test on all covariates across all data sets results in an average of 16.7% of the covariates being discovered, while for the $\ell$-test, this figure is 18.6% (this represents an 11% improvement), suggesting that the $\ell$-test’s theoretical benefits under sparsity provide genuine power gains in real data sets. Similarly, the $t$-test confidence intervals have an average width of 3.85 while the $\ell$-test confidence intervals’ is 3.54, representing about an 8% improvement. We also note that previous works studying the same data have aggregated discovered mutations to the gene level, and if we do this, the comparison remains similar: the $t$-test’s average power is 31.8% discovered genes while the $\ell$-test’s is 36.4% (a 14% improvement).
@@ 7 Discussion
The $\ell$-test leverages sparsity by using the LASSO coefficient estimate $|\hat{\beta}^{\lambda}_{j}|$ as its test statistic and can achieve power close to that of the one-sided $t$-test without any knowledge about the true sign of the coefficient. The $\ell$-test can be inverted to obtain confidence intervals that are about $10\%$ shorter than the $t$-test intervals in the sparse settings of our simulations, and the $\ell$-test and confidence intervals can also be adjusted for LASSO selection. The $\ell$-test can also be applied asymptotically to a rich class of models far beyond the homoskedastic Gaussian linear model.
The key idea underlying the $\ell$-test is that a belief about model structure can be leveraged via information in the null model’s sufficient statistic to improve (when that structure holds) inferential power. We believe this idea can be used for types of structure other than sparsity as well, e.g., if $\bm{\beta}$ is dense but smooth (i.e., it has small total variation), a test based on the fused LASSO (tibshirani2005sparsity) may be more appropriate. However, while the key idea of constructing a valid test based on a structure-leveraging test statistic as described at the beginning of Section 2 can always be carried out via co-sufficient sampling (CSS), the details in Sections 2.12.2 of what make the $\ell$-test computationally efficient and powerful rely heavily on the particulars of the LASSO, and hence we defer to future work the possibility of similar results for other types of structure.
@@ Appendix A Characterization of the $\ell$-distribution under $H_{j}(\gamma):\beta_{j}=\gamma$
Theorem 4.
Consider the linear model defined in Theorem 1, fix $\gamma\in\mathbb{R}$ and define $\hat{\sigma}_{j}(\gamma)=\|(\bm{I}-\bm{P}_{-j})(\bm{y}-\gamma\bm{X}_{j})\|$. Then,
  1. (a)
    $\bm{S}^{(j)}(\gamma)=(\bm{X}_{-j}^{T}\bm{y},\|(\bm{I}-\bm{P}_{-j})(\bm{y}-\gamma\bm{X}_{j})\|)$ is sufficient under $H_{j}(\gamma)$. Furthermore, fixing an orthogonal matrix $\bm{V}$ for the column-space of $\bm{X}_{-j}$ as in Lemma 1, there exists a unique vector $\bm{u}^{\gamma}\in\mathbb{S}^{n-d}$ such that $\bm{y}=\hat{\bm{y}}_{j}+\gamma(\bm{I}-\bm{P}_{-j})\bm{X}_{j}+\hat{\sigma}_{j}(\gamma)\bm{V}\bm{u}^{\gamma}$ and
@@
Note that, as discussed in Section 5, the proof of item 1 of the above theorem directly follows from Lemma 1 applied to $(\bm{y}-\gamma\bm{X}_{j},\bm{X})$. Analogous to the proof of Theorem 1 using Lemma 1, one can derive item 2 of Theorem 4 from its item 1, by defining $\bm{z}:=\bm{y}-\gamma\bm{X}_{j}$ and $\bm{\delta}$ by $\bm{\delta}_{-j}=\bm{\beta}_{-j}$ and $\delta_{j}=\beta_{j}-\gamma$ and copying exactly the same proof as in Appendix D.3 but by replacing $(\bm{y},\bm{\beta},\hat{\sigma}_{j})$ with $(\bm{z},\bm{\delta},\hat{\sigma}_{j}(\gamma))$ and then substituting back for $\bm{\delta}$ and $\bm{z}$ at the end. We thus skip an explicit proof of Theorem 4.
Here we would also like to draw attention to the fact that the function $f_{\bm{S}^{(j)}(\gamma)}^{\gamma}$, for any $\gamma\in\mathbb{R}$, is defined for any real number $u$ and not necessarily restricted to $[0,1]$ and also that $\Lambda_{j}(b,\mathrm{sign}(b))$ can take values outside the [1,1][-1,1] range. In fact if the values exceed this range, we can often draw conclusive insights about the behavior of the LASSO estimate $\hat{\beta}_{j}^{\lambda}$. For example, $\Lambda(0,\pm 1)<-1$ implies that $u_{1}^{\gamma}>\Lambda(0,1)$ and hence that for any $\bm{y}$ generated using the condition in Theorem 4 for a $\bm{u}^{\gamma}\in\mathbb{S}^{n-d}$, the resultant LASSO estimator of the $j^{\mathrm{th}}$ coefficient, $\hat{\beta}_{j}^{\lambda}$, will always be positive. This can happen in situations as we next described in Section B and can in-fact, result in the $\ell$-test producing exactly the one-sided $t$-test $p$-value.